OpenStudy (anonymous):

The ratio of quarters to nickels is 2 to 5. If the total value of all the coins is $4.50, how many coins are there?

OpenStudy (anonymous):

quarter=25 cents nickel=5cents let no: of quarters=x and no: of nickel =y x/y=2/5 no: of quarters, x=2y/5 25*(2y/5)+5y=450 15y=450 y=30 no: of nickel=30 no: of quarters, x=2y/5=(2*30)/5=12